# Beyond Pairwise: Empowering LLM Alignment With Ranked Choice Modeling

### v1

@@ 4.2 Results
AlpacaEval 2 Arena-Hard
Method LC (%) WR (%) WR (%)
Base Model 24.76 (0.42) 24.40 (1.44) 23.6 (21.9, 25.4)
CPO (Xu et al., 2024) 32.25 (0.31) 34.18 (1.59) 31.3 (29.5, 33.0)
IPO (Azar et al., 2024) 37.95 (0.33) 33.51 (1.63) 31.0 (29.2, 32.6)
ORPO (Hong et al., 2024) 35.01 (0.39) 28.72 (1.52) 27.5 (25.5, 29.3)
RRHF (Yuan et al., 2023) 31.04 (0.21) 25.36 (1.47) 27.1 (25.3, 28.9)
SLiC-HF (Zhao et al., 2023) 31.04 (0.21) 25.36 (1.47) 26.9 (25.2, 28.9)
KTO (Ethayarajh et al., 2024) 32.55 (0.35) 29.70 (1.54) 25.8 (24.0, 27.7)
DPO (Rafailov et al., 2023) 41.24 (0.35) 40.24 (1.66) 32.6 (30.6, 34.7)
R-DPO (Park et al., 2024) 39.88 (0.34) 38.01 (1.63) 32.3 (30.1, 34.3)
SimPO (Meng et al., 2024) 44.15 (0.25) 38.84 (1.58) 33.5 (31.3, 35.8)
Mallows-RMJ-PO-Pairwise 39.33 (0.28) 48.71 (1.67) 36.5 (34.3, 38.6)
MNL-PO-Discrete 41.33 (0.29) 48.08 (1.68) 35.6 (33.6, 37.4)
Mallows-RMJ-PO-Discrete 39.19 (0.28) 51.17 (1.67) 36.3 (34.6, 37.9)
MNL-PO-Top-2 40.12 (0.22) 47.69 (1.67) 35.8 (33.2, 38.2)
Mallows-RMJ-PO-Top-2 41.95 (0.26) 53.01 (1.68) 37.2 (35.0, 39.4)
  • Notes: The last eight rows correspond to preference optimization methods within the RCPO framework. Among these, the first three (DPO, R-DPO, SimPO) are from prior work, and the remaining five are new variants in this paper. Standard errors (in parentheses) follow AlpacaEval 2 metrics, and 95% confidence intervals (in parentheses) follow Arena-Hard WR.
@@
Alignment performance. All eight choice-based methods—three existing and five newly proposed RCPO variants—clearly outperform other preference optimization baselines across all evaluation metrics. Notably, the best-performing RCPO method, Mallows-RMJ-PO-Top-2, surpasses the strongest non-RCPO baseline, IPO, by 4.00 points on AlpacaEval LC, 19.5 points on AlpacaEval WR, and 6.2 points on Arena-Hard WR. These gains demonstrate the strong potential of choice modeling as a principled framework for aligning LLMs with human preferences.
Impact of feedback structure. Fixing the reward function as 2, we find that training on top-2 feedback generally leads to better performance than top-1. This reflects the benefit of richer feedback structure. On the other hand, we envision that there is an implicit trade-off between informativeness of feedback and the intrinsic noise/error in the data. Therefore, the performance may not be always increasing in the length of feedback. That is why we stop at top-2 before going even further, such as the full rankings.
@@
Impact of choice models. The selection of the choice model can have a substantial impact on performance. Fixing the reward function as 2, Mallows-RMJ performs strongly across feedback types. In the pairwise setting, it outperforms all baselines on AlpacaEval WR and Arena-Hard WR, even surpassing MNL models trained on richer top-2 feedback. Its performance further improves with discrete or top-2 choice data.
@@
Robustness check on Gemma-2-9B-it. As shown in Table 3, the Mallows-RMJ-PO-Top-2 achieves the best results on AlpacaEval WR and Arena-Hard WR while remaining competitive on LC, highlighting strong generalization across base models.
Table 3: Evaluation Results for Gemma-2-9B-it.
AlpacaEval 2 Arena-Hard
Method LC (%) WR (%) WR (%)
Base Model 46.74 (0.15) 31.81 (1.58) 43.3 (41.3, 45.7)
SimPO (Meng et al., 2024) 54.11 (0.17) 47.23 (1.68) 57.4 (55.3, 59.6)
DPO (Rafailov et al., 2023) 58.01 (0.18) 56.13 (1.66) 59.9 (57.6, 62.2)
Mallows-RMJ-PO-Top-2 55.64 (0.11) 59.82 (1.65) 60.9 (58.9, 62.6)
@@ Appendix A Related Work
DPO’s success has attracted significant research attention in the LLM alignment community, yielding numerous variants with alternative objectives. These include formulations based on ranking objectives (Yuan et al., 2023; Song et al., 2024; Liu et al., 2024), pairwise comparisons employing other preference models (Chen et al., 2025), other implicit reward formulations (Wang et al., 2023; Meng et al., 2024; Gupta et al., 2025), and approaches leveraging binary feedback on individual prompt-response pairs (Ethayarajh et al., 2024). Distinct from prior studies, we investigate novel forms of preference feedback that enable alignment methods to directly leverage richer choice-based signals.
@@
Social choice theory is a field of study that deals with the aggregation of individual preferences to form a collective decision. Current approaches to LLM alignment involves the experimental studies (Huang et al., 2024), conceptual frameworks (Prasad, 2018; Mishra, 2023; Dai & Fleisig, 2024; Conitzer et al., 2024; Zhi-Xuan et al., 2024), and axiomatic frameworks (Ge et al., 2024). In contrast to the standard social choice setting (Dai & Fleisig, 2024), where voters provide full rankings over all alternatives, we focus on aggregating ranked choice preferences.
@@ Appendix C Practical Scheme of Mallows-RMJ Methods
The objectives in 10, Theorem 4, and 14 cannot be directly optimized with gradient-based methods, since they involve step functions that are either discontinuous or flat almost everywhere, yielding zero gradients for most values of $\theta$. To make these objectives amenable to efficient optimization with preference data, we replace the indicator functions with a sigmoid approximation. This smoothing technique is widely adopted in the machine learning literature (see, e.g., (Qin et al., 2010; Bruch et al., 2019)). Specifically, we approximate $\mathbb{I}\{x<0\}$ using $\sigma(-\beta x)$, where $\beta$ serves as a hyperparameter that controls the tightness of the approximation. An illustration is provided in Figure 2(a).
@@ Appendix D Proofs
Proof of Theorem 1. The Multinomial Logit (MNL) model (McFadden, 1972) is one of the most widely used utility-based discrete choice models. It assumes that each alternative $y_{i}\in S$ is associated with a mean utility $u_{y_{i}}(x)=\nu_{y_{i}}(x)+\epsilon_{i}$, where $\nu_{y_{i}}(x)$ is a deterministic component and $\epsilon_{i}$ follows an independent Gumbel distribution. Under this assumption, the choice probability of selecting alternative $y_{w}$ from assortment $S$ takes the closed-form
$\mathbb{P}(y_{w}\mid S\ ;x)=\tfrac{e^{\nu_{y_{w}}(x)}}{\sum_{y_{i}\in S}e^{\nu_{y_{i}}(x)}}.$
@@
$\mathbb{P}_{\pi_{\theta}}(y_{w}\mid S\ ;x)=\tfrac{e^{\beta\log\tfrac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}}}{\sum_{y_{i}\in S}e^{\beta\log\tfrac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\mathrm{ref}}(y_{i}\mid x)}}}=\tfrac{1}{\sum_{y_{i}\in S}e^{\beta\log\tfrac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\mathrm{ref}}(y_{i}\mid x)}-\beta\log\tfrac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}}}.$
Maximizing the likelihood yields to the following objective:
$\min_{\pi_{\theta}}\;-\mathbb{E}_{(x,S,y_{w})\sim\mathcal{D}}\log\sigma\big(-\log\sum\nolimits_{y_{i}\in S\setminus\{y_{w}\}}\exp\big(\beta\log\tfrac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\mathrm{ref}}(y_{i}\mid x)}-\beta\log\tfrac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}\big)\big).$
$\square$
Proof of Theorem 2. Following Feng & Tang (2023), the probability that an individual selects a top-$k$ choice $\mu^{k}=y_{1}\succ y_{2}\succ\ldots\succ y_{k}$ out of an assortment $S\subseteq\mathcal{N}$ is
$\mathbb{P}(\mu^{k}\mid S\ ;x)=\prod_{i=1}^{k}\tfrac{e^{\nu_{y_{i}}(x)}}{\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{i-1}\}}e^{\nu_{y_{j}}(x)}}.$
By identifying the deterministic utility $\nu_{y}(x)$ as the reward function defined in 2, we have the optimal RLHF policy satisfies
$\mathbb{P}_{\pi_{\theta}}(\mu^{k}\mid S\ ;x)$ $=\prod_{i=1}^{k}\tfrac{e^{\beta\log\tfrac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\mathrm{ref}}(y_{i}\mid x)}}}{\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{i-1}\}}e^{\beta\log\tfrac{\pi_{\theta}(y_{j}\mid x)}{\pi_{\mathrm{ref}}(y_{j}\mid x)}}}$
$=\prod_{i=1}^{k}\tfrac{1}{\sum_{y_{j}\in S\setminus\{y_{1},\ldots,y_{i-1}\}}e^{\beta\log\tfrac{\pi_{\theta}(y_{j}\mid x)}{\pi_{\mathrm{ref}}(y_{j}\mid x)}-\beta\log\tfrac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\mathrm{ref}}(y_{i}\mid x)}}}.$
To maximize the likelihood estimation, our objective becomes:
$\square$
@@
Combining (9) and (18), we have the optimal RLHF policy $\pi_{\theta}(\cdot\mid x)$ for (17) satisfies

### v2

@@ 4.2 Main Results
AlpacaEval 2 Arena-Hard UltraFeedback
Method LC (%) WR (%) WR (%) WR (%)
Base Model 24.76 (0.42) 24.40 (1.44) 23.6 (21.9,25.4) 42.51 (1.04)
CPO (Xu et al., 2024) 32.25 (0.31) 34.18 (1.59) 31.3 (29.5,33.0) 58.71 (1.05)
IPO (Azar et al., 2024) 37.95 (0.33) 33.51 (1.63) 31.0 (29.2,32.6) 58.44 (1.05)
ORPO (Hong et al., 2024) 35.01 (0.39) 28.72 (1.52) 27.5 (25.5,29.3) 51.68 (1.06)
RRHF (Yuan et al., 2023) 31.04 (0.21) 25.36 (1.47) 27.1 (25.3,28.9) 44.28 (1.06)
SLiC-HF (Zhao et al., 2023) 31.04 (0.21) 25.36 (1.47) 26.9 (25.2,28.9) 44.09 (1.06)
KTO (Ethayarajh et al., 2024) 32.55 (0.35) 29.70 (1.54) 25.8 (24.0,27.7) 53.33 (1.06)
DPO (Rafailov et al., 2023) 41.24 (0.35) 40.24 (1.66) 32.6 (30.6,34.7) 62.36 (1.03)
R-DPO (Park et al., 2024) 39.88 (0.34) 38.01 (1.63) 32.3 (30.1,34.3) 60.68 (1.05)
SimPO (Meng et al., 2024) 44.15 (0.25) 38.84 (1.58) 33.5 (31.3,35.8) 50.17 (1.09)
DPO-AllPairs 33.02 (0.24) 38.47 (1.65) 29.6 (27.3,31.5) 51.95 (1.08)
Mallows-RMJ-PO-Pairwise 39.33 (0.28) 48.71 (1.67) 36.5 (34.3,38.6) 66.28 (1.02)
MNL-PO-Discrete 41.33 (0.29) 48.08 (1.68) 35.6 (33.6,37.4) 64.64 (1.03)
Mallows-RMJ-PO-Discrete 39.19 (0.28) 51.17 (1.67) 36.3 (34.6,37.9) 67.56 (1.01)
MNL-PO-Top-2 40.12 (0.22) 47.69 (1.67) 35.8 (33.2,38.2) 64.01 (1.04)
Mallows-RMJ-PO-Top-2 41.95 (0.26) 53.01 (1.68) 37.2 (35.0,39.4) 68.91 (1.00)
  • Notes: Methods are grouped by horizontal rules. The first block reports the base model. The second block reports existing preference optimization baselines. The third block reports prior preference optimization methods that can be viewed as special cases within the RCPO framework. DPO-AllPairs denotes the DPO method trained on pairwise data implied from the Top-2 ranking. The final block reports new preference optimization variants proposed in this paper under the RCPO framework. Standard errors (in parentheses) are reported for AlpacaEval 2 and UltraFeedback, and 95% CIs (in parentheses) are reported for Arena-Hard WR.
@@
Generalization performance. All nine choice-based methods—four existing and five newly proposed RCPO variants— outperform other preference optimization baselines across most evaluation metrics. Notably, the best-performing RCPO method, Mallows-RMJ-PO-Top-2, surpasses the strongest non-RCPO baseline, IPO, by 4.00 percentage points on AlpacaEval LC, 19.5 percentage points on AlpacaEval WR, 6.2 percentage points on Arena-Hard WR, and 9.47 percentage points on UltraFeedback WR. These gains demonstrate the strong potential of choice modeling as a principled framework for aligning LLMs with human preferences.
Impact of feedback structure. Fixing the reward function as 2, we find that training on top-2 feedback generally leads to better performance than top-1. This reflects the benefit of richer feedback structure. We will discuss the impact of longer feedback structures in the later ablation study.
@@
Impact of choice models. The selection of the choice model can have a substantial impact on performance. Fixing the reward function as 2, Mallows-RMJ performs strongly across feedback types. In the pairwise setting, it outperforms all baselines on AlpacaEval WR, Arena-Hard WR, and UltraFeedback WR, even surpassing MNL models trained on richer top-2 feedback. Its performance further improves with discrete or top-2 choice data. We also observe that, when trained on the same feedback structure, the two choice models require nearly identical training time. Further details can be found in the Appendix F.1.
@@
Robustness check on Gemma-2-9B-it and Mistral-7B-Instruct. As shown in Table 3, the Mallows-RMJ-PO-Top-2 consistently outperforms competitive baselines on the benchmarks, highlighting strong generalization across base models.
Table 3: Evaluation Results for Other Base Models.
Gemma-2-9B-it Mistral-7B-Instruct
AlpacaEval 2 Arena-Hard AlpacaEval 2 Arena-Hard
Method LC (%) WR (%) WR (%) LC (%) WR (%) WR (%)
Base Model 46.74 (0.15) 31.81 (1.58) 43.3 (41.3,45.7) 14.53 (0.44) 11.94 (1.09) 10.8 (9.5,12.0)
SimPO (Meng et al., 2024) 54.11 (0.17) 47.23 (1.68) 57.4 (55.3,59.6) 26.32 (0.38) 30.13 (1.54) 19.3 (17.8,20.6)
DPO (Rafailov et al., 2023) 58.01 (0.18) 56.13 (1.66) 59.9 (57.6,62.2) 23.32 (0.39) 19.91 (1.34) 16.8 (15.5,18.3)
Mallows-RMJ-PO-Top-2 55.64 (0.11) 59.82 (1.65) 60.9 (58.9,62.6) 29.57 (0.22) 37.58 (1.68) 16.9 (15.4,18.3)
@@
4.3 Ablation Study
The ablation studies are conducted using the Llama-3-8B-Instruct. Due to space constraints, the main text focuses on the impact of the ranked choice length $k$ and the assortment size $|S|$. The sample efficiency of training on ranked choice data, the effects of the $\beta$ parameter, and different treatments on ranked responses are deferred to the Appendix F.3.
4.3.1 Impact of ranked choice length $k$ and assortment size $|S|$.
We study the impact of the ranked choice length $k$ and assortment size $|S|$ from two perspectives: generalization performance and time efficiency. The results are summarized in Table 4 below.
Impact of $k$ on generalization performance. A large $k$ may not necessarily improve performance. For example, when $|S|=5$, $k=2$ is the best. One possible reason is that higher values of $k$ make the alignment process more reliant on high-quality data. Specifically, when $k$ is small, selecting the top-$k$ items is relatively easy, but as $k$ becomes larger, distinguishing the relative order among the remaining candidates may become more difficult. Hence a too large $k$ can introduce additional noise or place unnecessary burden on annotators. This finding is conceptually consistent with the prior literature in a best-item selection context (Feng and Tang, 2023). It also highlights the benefit of top-$k$ feedback compared to the more extreme options, such as pairwise comparisons or full ranking.
Impact of $|S|$ on generalization performance. Increasing $|S|$ from two to three and five generally leads to better performance. Notably, even $|S|=3$ can achieve significant improvement over pairwise ($|S|=2$). One possible reason is that incorporating more negative samples enables the language model to learn better distinctions.
Impact of $k$ and $|S|$ on time efficiency. The training time is insensitive to the value of $k$. It scales approximately linearly with the assortment size $|S|$. The reason is that when $|S|$ is larger, fewer responses can be contained in one batch, and therefore more time to complete an epoch.
In general, we believe that a moderately small $(|S|,k)$ will stand in the sweet spot between keeping feedback simple and making efficient use of data.
Table 4: Evaluation Results of Varied $k$ and $|S|$.
Method AlpacaEval 2 LC (%) AlpacaEval 2 WR (%) Arena-Hard WR (%) Training Time (Hours)
DPO 41.24 (0.35) 40.24 (1.66) 32.6 (30.6, 34.7) 1.5
Mallows-RMJ-PO-Pairwise 39.33 (0.28) 48.71 (1.67) 36.5 (34.3, 38.6) 1.5
assort3-Mallows-RMJ-PO-Discrete 40.28 (0.27) 50.49 (1.68) 36.3 (34.1, 38.8) 2.3
assort3-Mallows-RMJ-PO-Top-2 39.57 (0.27) 49.65 (1.68) 36.9 (34.7, 39.0) 2.3
assort5-Mallows-RMJ-PO-Discrete 39.19 (0.28) 51.17 (1.67) 36.3 (34.6, 37.9) 3.5
assort5-Mallows-RMJ-PO-Top-2 41.95 (0.26) 53.01 (1.68) 37.2 (35.0, 39.4) 3.5
assort5-Mallows-RMJ-PO-Top-3 41.05 (0.25) 52.59 (1.69) 37.7 (35.8, 39.8) 3.5
assort5-Mallows-RMJ-PO-Top-4 39.92 (0.26) 51.08 (1.68) 37.4 (35.5, 39.2) 3.5
  • Notes: Standard errors (in parentheses) follow AlpacaEval 2 metrics, and 95% confidence intervals (in parentheses) follow Arena-Hard WR.
@@ Appendix A Related Work
DPO’s success has attracted significant research attention in the LLM alignment community, yielding numerous variants with alternative objectives. These include formulations based on ranking objectives (Yuan et al., 2023; Song et al., 2024; Liu et al., 2024; Zhao et al., 2025; Cai et al., 2025), pairwise comparisons employing other preference models (Chen et al., 2025), multi-level preference data (Zhang et al., 2024), other implicit reward formulations (Wang et al., 2023; Meng et al., 2024; Gupta et al., 2025), and approaches leveraging binary feedback on individual prompt-response pairs (Ethayarajh et al., 2024). Distinct from prior studies, we investigate novel forms of preference feedback that enable alignment methods to directly leverage richer choice-based signals.
@@
Social choice theory is a field of study that deals with the aggregation of individual preferences to form a collective decision. Current approaches to LLM alignment involves the experimental studies (Huang et al., 2024), conceptual frameworks (Prasad, 2018; Mishra, 2023; Dai and Fleisig, 2024; Conitzer et al., 2024; Zhi-Xuan et al., 2024), and axiomatic frameworks (Ge et al., 2024). In contrast to the standard social choice setting (Dai and Fleisig, 2024), where voters provide full rankings over all alternatives, we focus on aggregating ranked choice preferences.
@@ Appendix C Practical Scheme of Mallows-RMJ Methods
The objectives in 10, Theorem 4, and 14 cannot be directly optimized with gradient-based methods, since they involve step functions that are either discontinuous or flat almost everywhere, yielding zero gradients for most values of $\theta$. To make these objectives amenable to efficient optimization with preference data, we replace the indicator functions with a sigmoid approximation. This smoothing technique is widely adopted in the machine learning literature (see, e.g., (Qin et al., 2010; Bruch et al., 2019)). Specifically, we approximate $\mathbb{I}\{x<0\}$ using $\sigma(-x)$. An illustration is provided in Figure 2(a).
@@ Appendix D Proofs
Proof of Theorem 1. The Multinomial Logit (MNL) model (McFadden, 1972) is one of the most widely used utility-based discrete choice models. It assumes that each alternative $y_{i}\in S$ is associated with a latent utility $u_{y_{i}}(x)=\nu_{y_{i}}(x)+\epsilon_{i}$, where $\nu_{y_{i}}(x)$ is a deterministic component and $\epsilon_{i}$ follows an independent Gumbel distribution. Under this assumption, the choice probability of selecting alternative $y_{w}$ from assortment $S$ takes the closed form
$\mathbb{P}(y_{w}\mid S\ ;x)=\tfrac{\exp(\nu_{y_{w}}(x))}{\sum_{y_{i}\in S}\exp(\nu_{y_{i}}(x))}.$
@@
$\mathbb{P}_{\pi_{\theta}}(y_{w}\mid S\ ;x)=\tfrac{e^{\beta\log\tfrac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}}}{\sum_{y_{i}\in S}e^{\beta\log\tfrac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\mathrm{ref}}(y_{i}\mid x)}}}.$
Maximizing the likelihood yields the following objective:
$\min_{\pi_{\theta}}\;-\mathbb{E}_{(x,S,y_{w})\sim\mathcal{D}}\left[\beta\log\tfrac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}-\log\sum_{y_{i}\in S}e^{\beta\log\tfrac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\mathrm{ref}}(y_{i}\mid x)}}\right],$
which is equivalent to the objective in 7. $\square$
Proof of Theorem 2. Following Feng and Tang (2023), the probability that an individual selects a top-$k$ choice $\mu^{k}=y_{1}\succ y_{2}\succ\ldots\succ y_{k}$ out of an assortment $S\subseteq\mathcal{N}$ is
$\mathbb{P}(\mu^{k}\mid S\ ;x)=\prod_{i=1}^{k}\tfrac{e^{\nu_{y_{i}}(x)}}{\sum_{j=i}^{k}e^{\nu_{y_{j}}(x)}+\sum_{y_{h}\in S\setminus\{y_{1},\ldots,y_{k}\}}e^{\nu_{y_{h}}(x)}}.$
By identifying the deterministic utility $\nu_{y}(x)$ as the reward function defined in 2, we have that the optimal RLHF policy satisfies
$\mathbb{P}_{\pi_{\theta}}(\mu^{k}\mid S\ ;x)=\prod_{i=1}^{k}\tfrac{e^{\beta\log\tfrac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\mathrm{ref}}(y_{i}\mid x)}}}{\sum_{j=i}^{k}e^{\beta\log\tfrac{\pi_{\theta}(y_{j}\mid x)}{\pi_{\mathrm{ref}}(y_{j}\mid x)}}+\sum_{y_{h}\in S\setminus\{y_{1},\ldots,y_{k}\}}e^{\beta\log\tfrac{\pi_{\theta}(y_{h}\mid x)}{\pi_{\mathrm{ref}}(y_{h}\mid x)}}}.$
To maximize the likelihood, our objective becomes:
$\min_{\pi_{\theta}}\;-\mathbb{E}_{(x,S,\mu^{k})\sim\mathcal{D}}\left[\sum_{i=1}^{k}\beta\log\tfrac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\mathrm{ref}}(y_{i}\mid x)}-\sum_{i=1}^{k}\log\Bigl(\sum_{j=i}^{k}e^{\beta\log\tfrac{\pi_{\theta}(y_{j}\mid x)}{\pi_{\mathrm{ref}}(y_{j}\mid x)}}+\sum_{\begin{subarray}{c}y_{h}\in\\ S\setminus\{y_{1},\ldots,y_{k}\}\end{subarray}}e^{\beta\log\tfrac{\pi_{\theta}(y_{h}\mid x)}{\pi_{\mathrm{ref}}(y_{h}\mid x)}}\Bigr)\right],$
which is equivalent to the objective in Theorem 2. $\square$
@@
Combining (9) and (18), we have that the optimal RLHF policy $\pi_{\theta}(\cdot\mid x)$ for (17) satisfies